# Bayesian variable selection in sample selection models using spike-and-slab priors

### v2

@@ Abstract
Sample selection models represent a common methodology for correcting bias induced by data missing not at random. It is well known that these models are not empirically identifiable without exclusion restrictions. In other words, some variables predictive of missingness do not affect the outcome model of interest. The drive to establish this requirement often leads to the inclusion of irrelevant variables in the model. A recent proposal uses adaptive LASSO to circumvent this problem, but its performance depends on the so-called covariance assumption, which can be violated in small to moderate samples. Additionally, there are no tools yet for post-selection inference for this model. To address these challenges, we propose two families of spike-and-slab priors to conduct Bayesian variable selection in sample selection models. These prior structures allow for constructing a Gibbs sampler with tractable conditionals, which is scalable to the dimensions of practical interest. We illustrate the performance of the proposed methodology through a simulation study and present a comparison against adaptive LASSO and stepwise selection. We also provide two applications using publicly available real data. An implementation and code to reproduce the results in this paper can be found at https://github.com/adam-iqbal/selection-spike-slab.
Keywords Gibbs sampling  $\cdot$ missing data  $\cdot$ sample selection  $\cdot$ spike-and-slab prior
@@ 1 Introduction
Sample selection occurs when a portion of the sample is missing non-randomly, leading to a sample that is not representative of the population of interest. This is prevalent in medical and social sciences, where the outcome of interest is non-randomly selected (Heckman, 1976, 1979; Wooldridge, 2010). As a result, the sample becomes biased, potentially distorting the findings and generalizability of the results. Heckman (1976, 1979) introduced a regression model to correct for sample selection bias in the case where the outcome variable is continuous. The main idea is to formulate a two-equation regression model, one for the outcome process and one for the selection process, allowing for correlated errors under the assumption of bivariate normality. This kind of models are known as sample selection models (or Heckman correction). Heckman (1979) proposed a two-stage estimator of the parameters of sample selection models, and established consistency and asymptotic normality of such estimator. After this influential development, several extensions have been proposed to alleviate challenges posed by the assumption of normality of the errors. These include van Hasselt (2011), who extended the outcome model to mixture-of-Gaussian error distributions. Marchenko and Genton (2012), who extended Heckman’s model by modelling the errors in the two-equation model using a bivariate-$t$ distribution, which accounts for heavy tails, coupled with maximum likelihood estimation of the parameters. Ogundimu and Hutton (2016) assumed bivariate skew-normal errors, which account for departures from symmetry. Other extensions include the analysis of binary outcomes subject to sample selection bias. One of the challenges in all of these models is that the user must specify the variables that enter the outcome and selection models. Entering the same variables in both the outcome and selection submodels often leads to collinearity problems. It has also been shown that even with “exclusion restriction” rules - a covariate included in the selection but not the outcome - collinearity issues may appear in two-step estimators (Leung and Yu, 2000). This challenge points to the need of formal tools for specifying the variables that enter the outcome and selection models.
@@
Recently, the use of variable selection methods in the context of sample selection models has attracted considerable attention. For instance, Ogundimu (2022) used an Adaptive LASSO penalty to select the variables that enter the outcome and selection models. The distributional regression sample selection modelling introduced by Wiemann et al. (2022) allows for the use of Bayesian variable selection priors (shrinkage and spike-and-slab priors). However, Wiemann et al. (2022) only provide an implementation of their proposed model using a Metropolis within Gibbs sampler, which is computationally expensive and requires careful calibration of the priors for the additive components. In a similar vein, Vera (2023) considered the use of shrinkage and spike-and-slab priors, but they restricted variable selection to only the selection equation, and their implementation similarly relies on a Metropolis within Gibbs sampler.
@@
We propose two families of continuous spike-and-slab priors to conduct Bayesian variable selection in sample selection models (George and McCulloch, 1993; Tadesse and Vannucci, 2021). The first class of priors allows for the use of scale mixtures of normals for the spike and slab components, which include Laplace and Student-$t$ distributions. The second one is defined conditionally on the marginal variance of the outcome equation, allowing for easier calibration, and also includes scale mixtures of normals for the spike and slab components. We provide a detailed discussion on the calibration of both priors, taking into account the different nature of the outcome and selection models. We derive a closed-form Gibbs sampler for the posterior distribution of the parameters, under both classes of priors, which scales well to the dimensions of interest in practical applications. Posterior samples can be used to jointly select variables for both the selection and outcome equations, without the need for an exclusion restriction, and to produce post-selection inference on the parameters.
@@ 2 Sample selection models
where $\sigma>0$ and $\rho\in(-1,1)$. Note that, in equation (8), the variance of the selection equation is set to $1$. This is because we only observe the sign $s_{i}$, so that this parameter cannot be identified. Additionally, the data is only partially observed. That is, we only observe the indicator $s_{i}$ and the selected sample $y_{i}$, where
$s_{i}=\mathbb{I}(s^{*}_{i}>0)\,,\quad y_{i}=s_{i}y^{*}_{i}\,.$
@@
$+$ $\sum_{\{s_{i}=1\}}\Bigg{(}\log\Phi\left(\frac{\alpha_{0}+\bm{w}_{i}^{\top}\bm{\alpha}+\rho(y_{i}-\beta_{0}-\bm{x}_{i}^{\top}\bm{\beta})/\sigma}{\sqrt{1-\rho^{2}}}\right)-\frac{1}{2}\left(\frac{y_{i}-\beta_{0}-\bm{x}_{i}^{\top}\bm{\beta}}{\sigma}\right)^{2}$
$\qquad\quad-\log(\sigma)-\frac{1}{2}\log(2\pi)\Bigg{)}\,.$
@@
In the Bayesian framework, van Hasselt (2011) showed that it is possible to construct a tractable Gibbs sampler for model (1) by using the reparametrization $\tilde{\sigma}^{2}=\sigma^{2}(1-\rho^{2})$ and $\tilde{\rho}=\rho\sigma$. The conditional distributions for $(s^{*}_{i},y^{*}_{i})$ under this reparametrization are
@@
where $\bm{\theta}=(\beta_{0},\bm{\beta}^{\top},\alpha_{0},\bm{\alpha}^{\top},\tilde{\rho},\tilde{\sigma}^{2})^{\top}$. Full details about the Gibbs sampler, and the corresponding priors, are presented in van Hasselt (2011) and reproduced in Appendix B for completeness. The major advantage of this Gibbs sampler is that every conditional distribution is in closed-form, allowing for fast simulations. The sampler does rely on $(p+1)\times(p+1)$ and $(q+1)\times(q+1)$ matrix inversions, but the dimensions of interest in applications for sample selection models are not so large that this is an onerous task. The Gibbs sampler also requires sampling from $n$ truncated normal random variables at each step, but this can be done relatively fast using available methods and R packages (Geweke, 1991). We will capitalize on van Hasselt (2011) to develop a Gibbs sampler in Section 4 for the priors proposed in Section 3.
@@ 3 Spike-and-slab prior formulation
Our aim is to perform Bayesian variable selection in sample selection models of type (1). To this end, let us define the variable inclusion indicators as follows. For $j=1,\ldots,p$, let $\gamma^{O}_{j}=1$ if $\beta_{j}$ is included in the outcome model, and $\gamma^{O}_{j}=0$ otherwise. Similarly, for $k=1,\ldots,q$, let $\gamma^{S}_{k}=1$ if $\alpha_{k}$ is included in the selection model, and $\gamma^{S}_{k}=0$ otherwise. Next, we present the two proposed classes of priors. The first class of priors contains continuous spike-and-slab priors with scale mixture of normal components (Tadesse and Vannucci, 2021). The second class of priors is defined conditionally on the marginal variance of the outcome model. Finally, we present a discussion on elicitation for each class of priors.
@@ 3.1 Priors
The Class I spike-and-slab prior is defined by the structure:
@@
with $\bm{v^{O}}$ and $\bm{v^{S}}$ defined as before. In sample selection models, the marginal variance for the outcome is $\tilde{\sigma}^{2}$, as shown in equation (10). Thus, the prior on the regression coefficients of the outcome model is defined conditionally on the marginal variance of the errors. This is a common strategy in the standard linear regression context (Louzada et al., 2023).
@@ 3.2 Prior calibration
We similarly choose $\tau_{1,\alpha}$ and $\tau_{1,\beta}$ in line with previous work on spike-and-slab priors. For the probit equation, we choose $\tau_{1,\alpha}=C_{\alpha}$, a constant. This is in line with the choice made by Narisetty et al. (2019) for logistic regression. We choose $C_{\alpha}=\sqrt{3}/\pi$ to match the probit scale. For the linear part of the model, we use $\tau_{1,\beta}=C_{\beta}(\log(n))^{1/2}$ similarly to Narisetty and He (2014). This ensures that as $n\rightarrow\infty$, the probability of including a non-zero effect will converge to $1$. The elicitation of $C_{\beta}$ is dependent on the context. For instance, we may choose it to include small effects as often as possible without losing specificity. However, this may be complicated for small sample sizes, such as $n=100$. Furthermore, the slab variance is less influential for large samples, so we elicitate based on a moderate sample size, such as $n=500$. Suppose that $\tau_{0,\beta}=1/\sqrt{n}$ and assume that $r=0.5$. For a small effect of magnitude $0.25$, there is variability in the posterior sampling, so we wish to ensure that we include marginally smaller effects, say of magnitude $0.15$, with high probability, but not effects significantly smaller. With $\tau_{1,\beta}=0.5$, the probability of including a coefficient sample of magnitude $0.15$ is approximately $96\%$. But an effect size of $0.05$ only has a $14\%$ probability of inclusion. As a result, we recommend the heuristic choice of $C_{\beta}=\left(4\log(500)\right)^{-1/2}$. It should be noted that while the above recommendations provide a reasonable trade-off between including small effects and excluding spurious ones, it may be desirable to focus on one over the other. For instance, if there is no interest in including small effects, one could use a larger value for $C_{\beta}$ such as $1$ to improve specificity.
@@
We do not directly define priors on $\rho$ and $\sigma^{2}$, but indirectly through $\tilde{\rho}$ and $\tilde{\sigma}^{2}$. The prior on $\tilde{\sigma}^{2}$ is $\text{IG}(c,d)$, with $c,d>0$ positive constants, and the prior on $\tilde{\rho}$ is $N(0,\tau\tilde{\sigma}^{2})$ where $\tau>0$ is a positive constant. We let the prior on $\tilde{\rho}$ depend on $\tilde{\sigma}^{2}$ so that the induced prior on $\rho$ has its density approach zero as $\rho\rightarrow\pm 1$. Choosing $c=d=1$ leads to an induced prior on $\sigma^{2}$ with similar behaviour to an Inverse Gamma. The choice of $\tau$ in the prior for $\tilde{\rho}$ has little impact on the posterior inference for $\bm{\gamma^{S}},\bm{\gamma^{O}},\alpha_{0},\beta_{0},\bm{\alpha},\bm{\beta}$ and $\sigma^{2}$. However, it does have an impact on the posterior inference of $\rho$. The induced prior on $\rho$ depends on the choice of $\tau$: for small values such as $\tau=0.5$, it behaves similarly to a uniform distribution, while for larger values like $\tau=5$, it behaves similarly to a $Beta(1/2,1/2)$ distribution. Empirically we found that $\tau=5$ leads to more accurate posterior means of $\rho$ than $\tau=0.5$. See Appendix A for a further discussion.
@@ 4 Gibbs samplers
The exact expressions for $\bm{a^{*}},\bm{A^{*}},\bm{b^{*}}$ and $\bm{B^{*}}$ can be found in Appendix B. Steps 1 through 7 have distributions that are well-known and can easily be sampled from. Steps 8, 9 and 10 depend on the choice of scale distributions $\pi_{0,\alpha},\pi_{1,\alpha},\pi_{0,\beta}$ and $\pi_{1,\beta}$. There is no closed form expression for a general scale mixing distribution, but certain choices lead to closed form sampling for these steps. A Laplace prior can be used by choosing the mixing distribution to be $\text{Exp}(1/2)$: in this case, the posterior distribution to sample from takes the form of an Inverse Gaussian (after conditioning in Step 8/9). A Student-$t$ prior can be used by choosing the mixing distribution to be Inverse Gamma, which is conjugate to the normal likelihood. In these cases and the normal case (by choosing the mixing distributions to be point masses at $1$), the sampler is in closed form.

### v3

@@ Abstract
Sample selection models represent a common methodology for correcting bias induced by data missing not at random. These models are not empirically identifiable without exclusion restrictions. In other words, some variables predictive of missingness do not affect the outcome model of interest. The drive to establish this requirement often leads to the inclusion of irrelevant variables in the model. A recent proposal uses adaptive LASSO to circumvent this problem, but its performance depends on the so-called covariance assumption, which can be violated in small to moderate samples. Additionally, there are no tools yet for post-selection inference for this model. To address these challenges, we propose two families of spike-and-slab priors to conduct Bayesian variable selection in sample selection models. These prior structures allow for constructing a Gibbs sampler with tractable conditionals, which is scalable to the dimensions of practical interest. We illustrate the performance of the proposed methodology through a simulation study and present a comparison against adaptive LASSO and stepwise selection. We also provide two applications using publicly available real data. An implementation and code to reproduce the results in this paper can be found at https://github.com/adam-iqbal/selection-spike-slab.
Keywords Gibbs sampling  $\cdot$ Heckman correction  $\cdot$ missing data  $\cdot$ prior elicitation  $\cdot$ scale mixtures
@@ 1 Introduction
Sample selection occurs when a portion of the sample is missing non-randomly, leading to a sample that is not representative of the population of interest. This is prevalent in medical and social sciences, where the outcome of interest is non-randomly selected (Heckman, 1976, 1979; Wooldridge, 2010). As a result, the sample becomes biased, potentially distorting the findings and generalizability of the results. Heckman (1976, 1979) introduced a regression model to correct for sample selection bias in the case where the outcome variable is continuous. The main idea is to formulate a two-equation regression model, one for the outcome process and one for the selection process, allowing for correlated errors under the assumption of bivariate normality. This kind of models are known as sample selection models (or Heckman correction). Heckman (1979) proposed a two-stage estimator of the parameters of sample selection models, and established consistency and asymptotic normality of such estimator. After this influential development, several extensions have been proposed to alleviate challenges posed by the assumption of normality of the errors. These include van Hasselt (2011), who extended the outcome model to mixture-of-Gaussian error distributions. Marchenko and Genton (2012), who extended Heckman’s model by modelling the errors in the two-equation model using a bivariate-$t$ distribution, which accounts for heavy tails, coupled with maximum likelihood estimation of the parameters. Ogundimu and Hutton (2016) assumed bivariate skew-normal errors, which account for departures from symmetry. Other extensions include the analysis of binary outcomes subject to sample selection bias. One of the challenges in these models is that the user must specify the variables that enter the outcome and selection models. Entering the same variables in both the outcome and selection submodels often leads to collinearity problems. It has also been shown that even with “exclusion restriction” rules - a covariate included in the selection but not the outcome - collinearity issues may appear in two-step estimators (Leung and Yu, 2000). This challenge points to the need for formal tools for specifying the variables that enter the outcome and selection models.
@@
Recently, the use of variable selection methods in the context of sample selection models has attracted considerable attention. For instance, Ogundimu (2022) used an Adaptive LASSO penalty to select the variables that enter the outcome and selection models. The distributional regression sample selection modelling introduced by Wiemann et al. (2022) allows for the use of Bayesian variable selection priors (shrinkage and spike-and-slab priors). However, Wiemann et al. (2022) only provide an implementation of their proposed model using a Metropolis within Gibbs sampler, which is computationally more expensive than a closed-form Gibbs sampler due to the need for iterative accept-reject Metropolis steps, and requires careful calibration of the priors for the additive components to ensure convergence. In a similar vein, Vera (2023) considered the use of shrinkage and spike-and-slab priors, but they restricted variable selection to only the selection equation, and their implementation similarly relies on a Metropolis within Gibbs sampler.
@@
We propose two families of continuous spike-and-slab priors (George and McCulloch, 1993; Tadesse and Vannucci, 2021) to conduct Bayesian variable selection in sample selection models. The first class of priors allows for the use of scale mixtures of normals for the spike and slab components, which include Laplace and Student-$t$ distributions. The second one is defined conditionally on the marginal variance of the outcome equation, allowing for easier calibration, and also includes scale mixtures of normals for the spike and slab components. We provide a detailed discussion on the calibration of both priors, taking into account the different nature of the outcome and selection models. We derive a closed-form Gibbs sampler for the posterior distribution of the parameters, under both classes of priors, which scales well to the dimensions of interest in practical applications. Posterior samples can be used to jointly select variables for both the selection and outcome equations, without the need for an exclusion restriction, and to produce post-selection inference on the parameters.
@@ 2 Sample selection models
where $\sigma>0$ and $\rho\in(-1,1)$. Note that, in equation (2), the variance of the selection equation is set to $1$. This is because we only observe the sign $s_{i}$, so that this parameter cannot be identified. Additionally, the data is only partially observed. That is, we only observe the indicator $s_{i}$ and the selected sample $y_{i}$, where $s_{i}=\mathbb{I}(s^{*}_{i}>0)$ and
$y_{i}=\begin{cases}y_{i}^{*}&\text{if }s_{i}=1\,,\\ \text{missing }&\text{if }s_{i}=0\,.\end{cases}$ (3)
@@
$+$ $\sum_{\{s_{i}=1\}}\Bigg{(}\log\Phi\left(\frac{\alpha_{0}+\bm{w}_{i}^{\top}\bm{\alpha}+\rho(y_{i}-\beta_{0}-\bm{x}_{i}^{\top}\bm{\beta})/\sigma}{\sqrt{1-\rho^{2}}}\right)$
$\qquad\quad-\frac{1}{2}\left(\frac{y_{i}-\beta_{0}-\bm{x}_{i}^{\top}\bm{\beta}}{\sigma}\right)^{2}-\log(\sigma)-\frac{1}{2}\log(2\pi)\Bigg{)}\,.$
@@
In the Bayesian framework, van Hasselt (2011) showed that it is possible to construct a tractable Gibbs sampler for the model defined by equations (1)-(3) by using the reparametrization $\tilde{\sigma}^{2}=\sigma^{2}(1-\rho^{2})$ and $\tilde{\rho}=\rho\sigma$. The conditional distributions for $(s^{*}_{i},y^{*}_{i})$ under this reparametrization are
@@
where $\bm{\theta}=(\beta_{0},\bm{\beta}^{\top},\alpha_{0},\bm{\alpha}^{\top},\tilde{\rho},\tilde{\sigma}^{2})^{\top}$. Full details about the Gibbs sampler, and the corresponding priors, are presented by van Hasselt (2011) and reproduced in Appendix C for completeness. The major advantage of this Gibbs sampler is that every conditional distribution is in closed-form, allowing for fast simulations. The sampler does rely on $(p+1)\times(p+1)$ and $(q+1)\times(q+1)$ matrix inversions but is rare for the problems of interest in sample selection literature to exhibit dimensions high enough for this to pose a serious concern. The Gibbs sampler also requires sampling from $n$ truncated normal random variables at each step, but this can be done relatively fast using available methods and R packages (Geweke, 1991). We will capitalize on van Hasselt (2011) to develop a Gibbs sampler in Section 4 for the priors proposed in Section 3.
@@ 3 Spike-and-slab prior formulation
Our aim is to perform Bayesian variable selection in sample selection models of type (1). To this end, let us define the variable inclusion indicators as follows. For $j=1,\ldots,p$, let $\gamma^{O}_{j}=1$ if $\beta_{j}$ is included in the outcome model, and $\gamma^{O}_{j}=0$ otherwise. Similarly, for $k=1,\ldots,q$, let $\gamma^{S}_{k}=1$ if $\alpha_{k}$ is included in the selection model, and $\gamma^{S}_{k}=0$ otherwise. Next, we present the two proposed classes of priors. The first class of priors contains continuous spike-and-slab priors with scale mixture of normal components (George and McCulloch, 1993). The second class of priors is defined conditionally on the marginal variance of the outcome model. Finally, we present a discussion on elicitation for each class of priors.
@@ 3.1 Priors
The Class I spike-and-slab prior (George and McCulloch, 1993) by the structure:
@@
with $\bm{v^{O}}$ and $\bm{v^{S}}$ defined as before. In sample selection models, the variance of the observed outcomes is $\tilde{\sigma}^{2}$, as shown in equation (5). Choosing the marginal variance $\sigma^{2}=\tilde{\sigma}^{2}+\tilde{\rho}^{2}$ would not lead to a natural expression for the posterior of $\tilde{\sigma}^{2}$, whereas $\tilde{\sigma}^{2}=\sigma^{2}(1-\rho^{2})$ does.
The advantage of Class II priors over Class I is that the magnitude of variables is only considered relative to the unexplained variance. This is a common strategy in the standard linear regression context (Louzada et al., 2023). The cost of using $\tilde{\sigma}^{2}$ instead of $\sigma^{2}$ is that the observed variance is less than the true variance, with $\tilde{\sigma}^{2}$ being shrunk dependent on $\rho$. For small magnitudes of $\rho$, the difference is small, but for large values, this can have a significant impact on the priors. For instance, when $\rho=0.7$, $\tilde{\sigma}^{2}=0.49\sigma^{2}$.
@@ 3.2 Prior calibration
We similarly choose $\tau_{1,\alpha}$ and $\tau_{1,\beta}$ in line with previous work on spike-and-slab priors. For the probit equation, we choose $\tau_{1,\alpha}=C_{\alpha}$, a constant. This is in line with the choice made by Narisetty et al. (2019) for logistic regression. We choose $C_{\alpha}=\sqrt{3}/\pi$ to match the probit scale. For the linear part of the model, we use $\tau_{1,\beta}=C_{\beta}(\log(n))^{1/2}$, which was used by Narisetty and He (2014) in their simulations. The discussion in Section 3.4 of Tadesse and Vannucci (2021) suggests that for this choice, as $n\rightarrow\infty$, the probability of including a non-zero effect will converge to $1$. This would be true for a fixed choice $\tau_{1,\beta}$ not dependent on $n$ as well, but the elicitation of $\tau_{1,\beta}$ is dependent on sample size. Even for moderately large sample sizes, a small choice of $\tau_{1,\beta}$ can shrink large coefficients and negatively affect the inference. But for small sample sizes, large $\tau_{1,\beta}$ will exclude small effects, losing sensitivity. Since this is not as much of an issue for moderately large sample sizes, we suggest a slab variance proportional to $\log(n)$.
@@
The elicitation of $C_{\beta}$ still depends on the context, and our default choice is derived as follows. Suppose that $\tau_{0,\beta}=1/\sqrt{n}$ and assume that $r=0.5$. For a small effect of magnitude $0.25$, there is variability in the posterior sampling, so we wish to ensure that we include marginally smaller effects, say of magnitude $0.15$, with high probability, but not effects significantly smaller. With $\tau_{1,\beta}=0.5$, the probability of including a coefficient sample of magnitude $0.15$ is approximately $96\%$. But an effect size of $0.05$ only has a $14\%$ probability of inclusion. As a result, we recommend the heuristic choice of $C_{\beta}=\left(4\log(500)\right)^{-1/2}$. It should be noted that while the above recommendations provide a reasonable trade-off between including small effects and excluding spurious ones, it may be desirable to focus on one over the other. For instance, if there is no interest in including small effects, one could use a larger value for $C_{\beta}$ such as $1$ to improve specificity.
@@
We do not directly define priors on $\rho$ and $\sigma^{2}$, but indirectly through $\tilde{\rho}$ and $\tilde{\sigma}^{2}$. The prior on $\tilde{\sigma}^{2}$ is $\text{IG}(c,d)$, with $c,d>0$ positive constants, and the prior on $\tilde{\rho}$ is $N(0,\tau\tilde{\sigma}^{2})$ where $\tau>0$ is a positive constant. We let the prior on $\tilde{\rho}$ depend on $\tilde{\sigma}^{2}$ so that the induced prior on $\rho$ has its density approach zero as $\rho\rightarrow\pm 1$. Choosing $c=d=1$ leads to an induced prior on $\sigma^{2}$ with similar behaviour to an Inverse Gamma. The choice of $\tau$ in the prior for $\tilde{\rho}$ has little impact on the posterior inference for $\bm{\gamma^{S}},\bm{\gamma^{O}},\alpha_{0},\beta_{0},\bm{\alpha},\bm{\beta}$ and $\sigma^{2}$. However, it does have an impact on the posterior inference of $\rho$. The induced prior on $\rho$ depends on the choice of $\tau$: for small values such as $\tau=0.5$, it behaves similarly to a uniform distribution, while for larger values like $\tau=5$, it behaves similarly to a $Beta(1/2,1/2)$ distribution. Empirically we found that $\tau=5$ leads to more accurate posterior means of $\rho$ than $\tau=0.5$. See Appendix B for a further discussion.
@@ 4 Gibbs samplers
The exact expressions for $\bm{a^{*}},\bm{A^{*}},\bm{b^{*}}$ and $\bm{B^{*}}$ can be found in Appendix C. Steps 1 through 7 have distributions that are well-known and can easily be sampled from. Steps 8, 9 and 10 depend on the choice of scale distributions $\pi_{0,\alpha},\pi_{1,\alpha},\pi_{0,\beta}$ and $\pi_{1,\beta}$. There is no closed form expression for a general scale mixing distribution, but certain choices lead to closed form sampling for these steps. A Laplace prior can be used by choosing the mixing distribution to be $\text{Exp}(1/2)$: in this case, the posterior distribution to sample from takes the form of an Inverse Gaussian (after conditioning in Step 8/9). A Student-$t$ prior can be used by choosing the mixing distribution to be Inverse Gamma, which is conjugate to the normal likelihood. In these cases and the normal case (by choosing the mixing distributions to be point masses at $1$), the sampler is in closed form.